# MIMO: Controllable Character Video Synthesis with Spatial Decomposed Modeling

### v1

@@ Abstract
Character video synthesis aims to produce realistic videos of animatable characters within lifelike scenes. As a fundamental problem in the computer vision and graphics community, 3D works typically require multi-view captures for per-case training, which severely limits their applicability of modeling arbitrary characters in a short time. Recent 2D methods break this limitation via pre-trained diffusion models, but they struggle for pose generality and scene interaction. To this end, we propose MIMO, a novel framework which can not only synthesize character videos with controllable attributes (i.e., character, motion and scene) provided by simple user inputs, but also simultaneously achieve advanced scalability to arbitrary characters, generality to novel 3D motions, and applicability to interactive real-world scenes in a unified framework. The core idea is to encode the 2D video to compact spatial codes, considering the inherent 3D nature of video occurrence. Concretely, we lift the 2D frame pixels into 3D using monocular depth estimators, and decompose the video clip to three spatial components (i.e., main human, underlying scene, and floating occlusion) in hierarchical layers based on the 3D depth. These components are further encoded to canonical identity code, structured motion code and full scene code, which are utilized as control signals of synthesis process. The design of spatial decomposed modeling enables flexible user control, complex motion expression, as well as 3D-aware synthesis for scene interactions. Experimental results demonstrate the proposed method’s effectiveness and robustness.
Figure 1: Given a single reference character, MIMO can synthesize animated avatars in driving 3D poses retrieved from motion datasets (left) or extracted from in-the-wild videos (right). Real-world scenes from driving videos can also be integrated into the synthesis with natural human-object interactions. MIMO simultaneously achieves advanced scalability to arbitrary characters, generality to novel 3D motions, and applicability to interactive real-world scenes in a unified framework.
@@ 1 Introduction
The aim of this paper is to propose a brand-new and boosting method for controllable video synthesis, which can not only synthesize character videos with controllable attributes (i.e., character, motion and scene) provided by very simple user inputs, but also achieve advanced scalability to arbitrary characters, generality to novel 3D motions, and applicability to interactive real-world scenes in a unified framework (see Figure 1). In other words, the proposed method is capable of mimicking anyone anywhere with complex motions and object interactions, thus named MIMO. As more concretely illustrated in Figure 2, users are allowed to feed multiple inputs (e.g., a single image for character, a pose sequence for motion, and a single video/image for scene) to provide desired attributes respectively or a direct driving video as input. The proposed model can embed target attributes into the latent space to construct target codes or encode the driving video with spatial-aware decomposition as spatial codes, thus enabling intuitive attribute control of the synthesis by freely integrating latent codes in a specific order.
Our task setting significantly decreases the cost of video creation and enables wide applications for not only character animation, but also video attribute editing (e.g., character replacement, motion transfer and scene insertion). However, it is extremely challenging due to the simplicity of user inputs, the complexity of real-world scenarios and the absence of annotation for 2D videos. With the great progress of 3D neural representations (e.g., NeRF [22] and 3D Gaussian splatting [12]), a series of works [23, 17, 27, 8, 15] tend to represent the dynamic human as a pose-conditioned NeRF or Gaussian to learn animatable avatars in high-fidelity rendering quality. However, they typically require fitting a neural field to multi-view captures or a monocular video of dynamic performers, which severely limits their applicability due to inefficient training and expensive data acquisition. Another 3D works explored faster and cheaper solutions by directly inferring 3D models from single human images, following by rigged animation and physical rendering [10, 9, 16, 21]. Unfortunately, the realism of the renderings is marginally compromised due to cumulative errors in sequential processes. Recently, several efforts [7, 30, 37, 26] have investigated the potential of 2D diffusion models on image guided character video synthesis. They show that high-fidelity character synthesis can be achieved by inserting image feature via a reference-net [7, 37] or control-net [30, 32] to a pretrained diffusion model. However, they only focus on character animation in simple 2D motions (e.g., frontal dancing) and are less effective for articulated human motion in 3D space with limited pose generality. Moreover, they fail to produce lifelike video with complicated scenes accompanied by human-object interactions. We argue that the cause for these difficulties stems from insufficient video attribute parser considered only in 2D feature space, thereby disregarding the inherent 3D nature of video occurrence.
@@
Figure 2: The basic idea of MIMO. Controllable character video synthesis with desired attributes provided by multiple inputs (e.g., a single image for character, a pose sequence for motion, and a single video/image for scene) or a driving video. Target attributes are embedded into the latent space as the target codes and the driving video is spatially decomposed as the spatial codes. Target character videos can be generated in user control with the combined attribute codes.
@@
To tackle these challenges, we propose a novel framework for controllable character video synthesis via spatial decomposed modeling. The core idea is to decompose and encode the 2D video in 3D-aware manner and employ more adequate expressions (e.g., 3D representations) for articulated properties. In contrast to previous works [7, 34] directly learn the whole 2D feature at each video frame, we lift the 2D frame pixels into 3D, and construct the decomposed spatial representations in 3D space, which are equipped with richer contextual information and can be used for control signals of synthesis process. Specifically, we decompose the video clip to three spatial components (scene, human and occlusion) in hierarchical layers based on 3D depth. In particular, human represents the main object in the video, scene represents the underlying background, and occlusion traces floating foreground objects. For the human component, we further disentangle the identity property via canonical appearance transfer and encode the 3D motion representation via structural body codes. The scene and occlusion components are embedded with a shared VAE encoder and re-organized as a full scene code. The decomposed latent codes are inserted as conditions of a diffusion-based decoder to reconstruction the video clip. In this way, the network learns not only controllable synthesis of various attributes, but also 3D-aware composition of main object, foreground and background. Thereby, it enables flexible user controls as well as challenging cases of complicated 3D motions and natural object interactions. In summary, our contributions are threefold:
  • We present a new approach capable of synthesizing realistic character videos with controllable attributes by directly providing simple user inputs, and simultaneously achieving advanced scalability, generality and applicability in a unified framework.
  • We propose the Spatial Decomposed Diffusion, a novel video generative model with automatic separation of spatial attributes, which enables not only flexible user control, but also 3D-aware synthesis for scene interactions.
  • We tackle the challenge of inadequate pose representation for articulated human by introducing structured body codes to express complex motions in spatial space, making an advanced generality to novel 3D motions.
@@ 2 Method Description
Our goal is to synthesize high-quality character videos with user-controlled visual attributes, such as character, motion and scenes. The desired attributes can be automatically extracted from an in-the-wild character video or simply provided by a single image, a pose sequence, and a single video/image, respectively. Different from previous methods using only weak control signals (e.g., text prompt) [35, 19] or insufficient 2D expressions [7, 34], our model achieves automatic and unsupervised separation of spatial components and encodes them into compact latent codes considering inherent 3D nature to control the synthesis. Thus, our dataset can only contain 2D character videos $\{v\in\mathbb{R}^{N\times H\times W}\}$ without any annotations.
The overview of the proposed framework is illustrated in Figure 3. Given a video clip $v$, MIMO learns a reconstruction process with automatic attribute encoding and composed condition decoding. Considering 3D nature of video occurrence, we extract the three spatial components in hierarchical layers based on 3D depth (Section 2.1). The first component of human is encoded with disentangled properties of identity and motion (Section 2.2). The last two components of scene and occlusion are embedded with a shared encoder and re-organized as a latent code (Section 2.3). These latent codes $\mathcal{C}$ are inserted into a diffusion decoder $\mathcal{D}$ as composed conditions (Section 2.4). $\mathcal{C}$, $\mathcal{D}$ are jointly learned by minimizing the difference between the synthesized frames and input frames in noise level (Section 2.5).
@@ 2.2 Disentangled human encoding
Structured motion. We define a set of latent codes $\mathcal{Z}=\{z_{1},z_{2},…,z_{nv}\}$, and anchor them to corresponding vertices of a deformable human body model (SMPL) [18], where $nv$ is the number of vertices. For the frame $t$, SMPL parameters $\mathcal{S}_{t}$ and camera parameters $\mathcal{C}_{t}$ are estimated from the monocular video frame $v^{h}_{t}$ using [3]. The spatial locations of the latent codes are then transformed based on the human pose $\mathcal{S}_{t}$ and projected to the 2D plane based on the camera setting $\mathcal{C}_{t}$. Using a differentiable rasterizer [14] with vertex interpolation, the 2D feature map $\mathcal{F}_{t}$ in continuous values can be obtained. $\{\mathcal{F}_{t},t=1,..,N\}$ will be stacked along the time axis and embedded into the latent space as the motion code $\mathcal{C}_{mo}$ by a pose encoder. In this way, we establish a correspondence that maps the same set of latent codes of underlying 3D body surface to the posed 2D renderings at different frames of arbitrary videos. This structured body codes enables more dense pose representation with 3D occlusions.
Canonical identity. To fully disentangle the appearance from posed video frames, an ideal solution is to learn the dynamic human representation from the monocular video and transform it from the posed space to the canonical space. Considering the efficiency, we employ a simplified method that directly transforms the posed human image to the canonical result in standard A-pose using a pretrained human repose model. The synthesized canonical appearance image is fed to ID encoders to obtain the identity code $\mathcal{C}_{id}$. This simple design enables full disentanglement of identity and motion attributes. Following [7], the ID encoders include a CLIP image encoder and a reference-net architecture to embed for the global and local feature, respectively, which compose $\mathcal{C}_{id}$.
@@ 2.3 Scene and occlusion encoding
In scene and occlusion branches, we use a shared and fixed VAE encoder [13] to embed the $v^{s}$ and $v^{o}$ into the latent space as the scene code $\mathcal{C}_{s}$ and occlusion code $\mathcal{C}_{o}$, respectively. Before $v^{s}$ input, we pre-recover it by a video inpainting method [36] for $\mathcal{R}(v^{s})$ to avoid the confusion brought by mask contours. Then the scene code $\mathcal{C}_{s}$ and the occlusion code $\mathcal{C}_{o}$ are concatenated together in order to get the full scene code $\mathcal{C}_{so}$ for composed synthesis. The independent encoding of spatial components (i.e., middle human, underlying scene, and floating occlusion) enable the network to learn an automatic layer composition, thus achieving natural character insertion in complicated scenes even with occluded object interactions.
@@ 2.4 Composed decoding
Given the latent codes of decomposed attributes, we re-compose them as conditions of the diffusion-based decoder for video reconstruction. As shown in Figure 4, we adapt denoising U-Net backbone built upon Stable Diffusion (SD) [25] with temporal layers from [4]. The full scene code $\mathcal{C}_{so}$ is concatenated with the latent noise, and is fed into a 3D convolution layer for fusion and alignment. The motion code $\mathcal{C}_{mo}$ is added to the fused feature and input to the denoising U-Net. For identity code $\mathcal{C}_{id}$, its local feature and global feature are inserted into the U-Net via self-attention layers and cross-attention layers, respectively. Finally, the denoised result is converted into the video clip $\hat{v}$ via a pretrained VAE decoder [13].
@@ 2.5 Training
We initialize the model of denoising U-Net and reference-net based on the pretrained weights from SD 1.5 [25], whereas the motion module is initialized with the weights of AnimateDiff [4]. During training, the weights of VAE encoder and decoder, as well as the CLIP image encoder are frozen. We optimize the denoising U-Net, pose encoder and reference-net with the diffusion noise-prediction loss:
@@
where $x_{0}$ is the augmented input sample, $t$ denotes the diffusion timestep, $x_{t}$ is the noised sample at $t$, and $\epsilon_{\theta}$ represents the function of the denoising UNet. We conduct the training on 8 NVIDIA A100 GPUs. It takes around 50k iterations with 24 video frames and a batch size of 4 for converge.

### v2

@@ Abstract
Character video synthesis aims to produce realistic videos of animatable characters within lifelike scenes. As a fundamental problem in the computer vision and graphics community, 3D works typically require multi-view captures for per-case training, which severely limits their applicability of modeling arbitrary characters in a short time. Recent 2D methods break this limitation via pre-trained diffusion models, but they struggle for flexible controls, pose generality and scene interaction. To this end, we propose MIMO, a novel framework which can not only synthesize realistic character videos with controllable attributes (i.e., character, motion and scene) provided by simple user inputs, but also simultaneously achieve advanced scalability to arbitrary characters, generality to novel 3D motions, and applicability to interactive real-world scenes in a unified framework. The core idea is to encode the 2D video to compact spatial codes, considering the inherent 3D nature of video occurrence. Concretely, we lift the 2D frame pixels into 3D using monocular depth estimators, and decompose the video clip into three spatial components (i.e., main human, underlying scene, and floating occlusion) in hierarchical layers based on the 3D depth. These components are further encoded to canonical identity code, structured motion code and full scene code, which are utilized as control signals of the synthesis process. The design of spatial decomposed modeling enables flexible user control, complex motion expression, as well as 3D-aware synthesis for scene interactions. Experimental results show that the proposed method outperforms prior works by a large margin in character animation synthesis and is effective in providing a high degree of controllability (i.e., arbitrary characters, novel 3D motions, interactive scenes), thus enabling brand-new editing tasks (e.g., video character replacement).
Figure 1: Given a single reference image of character, MIMO can synthesize animated avatars in driving 3D poses (visualized as skeleton sequences) retrieved from motion datasets (left) or extracted from in-the-wild videos (right). Real-world scenes from driving videos can also be integrated into the synthesis with natural human-object interactions. MIMO simultaneously achieves advanced scalability to arbitrary characters, generality to novel 3D motions, and applicability to interactive real-world scenes in a unified framework.
@@ 1 Introduction
The aim of this paper is to propose a brand-new and boosting method for controllable video synthesis, which can not only synthesize character videos with controllable attributes (i.e., character, motion and scene) provided by very simple user inputs, but also achieve advanced scalability to arbitrary characters, generality to novel 3D motions, and applicability to interactive real-world scenes in a unified framework (see Figure 1). In other words, the proposed method is capable of mimicking anyone anywhere with complex motions and object interactions, thus named MIMO. As more concretely illustrated in Figure 2, users are allowed to feed multiple inputs (e.g., a single image for character, a pose sequence for motion, and a single video even an image for scene) to provide desired attributes respectively or a direct driving video as input. The proposed model can embed target attributes into the latent space to construct target codes or encode the driving video with spatial-aware decomposition as spatial codes, thus enabling intuitive attribute control of the synthesis by freely integrating latent codes in a specific order.
@@
Figure 2: The basic idea of MIMO. Controllable character video synthesis with desired attributes provided by multiple inputs (e.g., a single image for character, a pose sequence for motion, and a single video even an image for scene) or a driving video. Target attributes are embedded into the latent space as the target codes and the driving video is spatially decomposed as the spatial codes. Target character videos can be generated in user control with the combined attribute codes.
Our task setting significantly decreases the cost of video creation and enables wide applications for not only character animation, but also video attribute editing (e.g., character replacement, motion transfer and scene insertion). However, it is extremely challenging due to the simplicity of user inputs, the complexity of real-world scenarios and the absence of 2D video annotations. With the great progress of 3D neural representations (e.g., NeRF [30] and 3D Gaussian splatting [19]), a series of works [31, 25, 41, 13, 22] tend to represent the dynamic human as a pose-conditioned NeRF or Gaussian to learn animatable avatars in high-fidelity rendering quality. However, they typically require fitting a neural field to multi-view captures or a monocular video of dynamic performers, which severely limits their applicability due to inefficient training and expensive data acquisition. Another 3D works explored faster and cheaper solutions by directly inferring 3D models from single human images, following by rigged animation and physical rendering [15, 14, 23, 29]. Unfortunately, the realism of the renderings is marginally compromised due to cumulative errors in sequential processes. Recently, several efforts [12, 44, 53, 39] have investigated the potential of 2D diffusion models on image-guided character video synthesis, named character animation. They show that high-fidelity character video can be synthesized by inserting image feature extracted from the reference-net [12, 53] or control-net [44, 47] into a pretrained diffusion model. However, they only focus on character synthesis in simple 2D motions (e.g., frontal dancing) and are less effective for articulated human motion in 3D space due to limited pose generality. Moreover, they fail to produce lifelike video for complicated scenes accompanied by human-object interactions and large camera movements. We argue that the cause for these difficulties stems from insufficient video attribute parser considered only in 2D feature space, thereby disregarding the inherent 3D nature of video occurrence.
To tackle these challenges, we propose a novel framework for controllable character video synthesis via spatial decomposed modeling. The core idea is to decompose and encode the 2D video in 3D-aware manner and employ more adequate expressions (e.g., 3D representations) for articulated properties. In contrast to previous works [12, 50] directly learn the whole 2D feature of each video frame, we lift the 2D frame pixels into 3D, and construct the decomposed spatial representations in 3D space, which are equipped with richer contextual information and can be used for control signals of the synthesis process. Specifically, we decompose the video clip to three spatial components (scene, human and occlusion) in hierarchical layers based on 3D depth. In particular, human represents the main object in the video, scene represents the underlying background, and occlusion traces floating foreground objects. For the human component, we further disentangle the identity property via canonical appearance transfer and encode the 3D motion representation via structural body codes. The scene and occlusion components are embedded with a shared VAE encoder and re-organized as a full scene code. The decomposed latent codes are inserted as conditions of a diffusion-based decoder to reconstruction the video clip. In this way, the network learns not only controllable synthesis of various attributes, but also 3D-aware layer composition of main object, foreground and background. Thereby, it enables flexible user controls as well as challenging cases of complicated 3D motions and natural object interactions. In summary, our contributions are threefold:
  • We propose a brand-new task that synthesizes character videos with controllable attributes by directly providing simple user inputs, and solve it with a novel approach to simultaneously achieve advanced scalability to arbitrary characters, generality to novel 3D motions, and applicability to interactive scenes in a unified framework.
  • We introduce the spatial decomposed modeling, an effective architecture to simulate intricate video observations by encoding the inherent spatial components. It enables not only flexible user control, but also 3D-aware synthesis in human-object interaction contexts.
  • We tackle the challenge of inadequate pose representation for articulated human by introducing structured motion codes. It provides better expressive ability to handle complicated motions in spatial space, thus enabling advanced generality of generative model to novel 3D motions.
@@
2 Related Work
3D Human Modeling. Since the introduction of Neural Radiance Fields (NeRF) [30], neural human representation has achieved remarkable success in obtaining articulated human models by fitting implicit neural fields to multi-view captures [32, 25, 36] or a monocular video [41, 17, 16]. HumanNeRF [41] proposes to represent human from a single video of a moving person by optimizing canonical volume and motion fields. NeuMan [17] jointly learns the decomposition of the human and the scene capable of novel pose rendering and animation of human in the scene. HOSNERF [24] extend to support human-environment interactions by introduce a dynamic human-object model. Recent works [22, 13] further introduce 3D Gaussian splatting [19] for realistic renderings. Despite achieving promising results, these methods typically require expensive data acquisition with precise camera/pose estimates and are less efficient for training and rendering, severely limiting their real-world applications. Another series of work explored faster and cheaper human modeling solutions via generative models  [9, 46, 3, 29]. En3D [29] learns an enhanced 3D human generator with efficient refiner to directly infer 3D models from single images or text prompts in few minutes. These generated avatars are rendered in a canonical body pose and aligned to an underlying 3D skeleton which allows for easy animation and the generation of motion videos. However, the fidelity of the driven results is marginally compromised due to cumulative errors inherent in the rendering processes.
Diffusion-based Character Video Synthesis. The remarkable progress in diffusion models has demonstrated promising results in image and video generation [7, 35, 8, 10, 2]. A plethora of methodologies proposed to incorporate these pre-trained diffusion models for human-centric video synthesis [38, 27, 4, 12, 44, 53, 50], enabling the transformation of character images into animated videos controlled by desired pose sequences. MagicAnimate [44] utilizes ControlNet [47] and an appearance encoder for identity preservation and pose guidance, building upon a video diffusion model. Animate Anyone [12] employs a UNet-based ReferenceNet to extract detailed features from reference images. Champ [53] further introduces a 3D parametric model to extract motion guidance as conditions. MimicMotion [50] presents a confidence-aware pose guidance approach to enhance generation quality and temporal smoothness. Despite producing visually appealing results, these methods encounter quality degradation issues in complex motion scenarios and are incapable of handling occlusion-aware generation in human-object interaction contexts. Our method built on diffusion models overcomes these challenges by a novel generative architecture with spatial decomposed modeling, considering inherent 3D nature for 2D videos.
@@ 3 Method Description
Our goal is to synthesize high-quality character videos with user-controlled visual attributes, such as character, motion and scenes. The desired attributes can be automatically extracted from an in-the-wild character video or simply provided by a single image, a pose sequence, and a single video, respectively. Different from previous methods using only weak control signals (e.g., text prompt) [51, 27] or insufficient 2D expressions [12, 50], our model achieves automatic and unsupervised separation of spatial components and encodes them into compact latent codes considering inherent 3D nature to control the synthesis. Thus, our dataset can only contain 2D character videos $\{v\in\mathbb{R}^{N\times H\times W}\}$ without any annotations.
The overview of the proposed framework is illustrated in Figure 3. Given a video clip $v$, MIMO learns a reconstruction process with automatic attribute encoding and composed condition decoding. Considering 3D nature of video occurrence, we extract the three spatial components in hierarchical layers based on 3D depth (Section 3.1). The first component of human is encoded with disentangled properties of identity and motion (Section 3.2). The last two components of scene and occlusion are embedded with a shared encoder and re-organized as a scene code (Section 3.3). These latent codes $\mathcal{C}$ are inserted into a diffusion decoder $\mathcal{D}$ as composed conditions (Section 3.4). $\mathcal{C}$, $\mathcal{D}$ are jointly learned by minimizing the difference between the generated frames and input frames via noise prediction (Section 3.5).
@@ 3.2 Disentangled human encoding
Structured motion code. We define a set of latent codes $\mathcal{Z}=\{z_{1},z_{2},…,z_{6890}\}$, and anchor them to corresponding vertices of a deformable human body model (SMPL) [26]. For the frame $t$, SMPL parameters $\mathcal{S}_{t}$ and camera parameters $\mathcal{C}_{t}$ are estimated from the monocular video frame $v^{h}_{t}$ using [5]. The spatial locations of the latent codes are then transformed based on the human pose $\mathcal{S}_{t}$ and projected to the 2D plane based on the camera setting $\mathcal{C}_{t}$. Using a differentiable rasterizer [21] with vertex interpolation, the 2D feature map $\mathcal{F}_{t}$ in continuous values can be obtained. $\{\mathcal{F}_{t},t=1,..,N\}$ will be stacked along the time axis and embedded into the latent space as the motion code $\mathcal{C}_{mo}$ by a pose encoder $\mathcal{E}_{p}$. In this way, we establish correspondences between the same set of identifiable latent codes on underlying 3D body surface and posed 2D renderings at different frames of arbitrary videos. This structured motion code enables clearer and more dense pose representation for articulated 3D motions in spatial space.
Canonical appearance transfer. To fully disentangle the appearance from posed video frames, an ideal solution is to learn the dynamic human representation from the monocular video and transform it from the posed space to the canonical space. Considering the efficiency, we employ a simplified method that directly transforms the posed human image to the canonical result in standard A-pose using a pretrained human repose model. The synthesized canonical appearance image is fed to ID encoders to obtain the identity code $\mathcal{C}_{id}$. This simple design enables full disentanglement of identity and motion attributes. Following [12], the ID encoders include a CLIP image encoder and a reference-net architecture to embed for the global and local feature, respectively, which composes $\mathcal{C}_{id}$.
@@ 3.3 Scene and occlusion encoding
In scene and occlusion branches, we use a shared and fixed VAE encoder [20] to embed the $v^{s}$ and $v^{o}$ into the latent space as the scene code $\mathcal{C}_{s}$ and occlusion code $\mathcal{C}_{o}$, respectively. Before $v^{s}$ input, we pre-recover it by a video inpainting method [52] as $\mathcal{R}(v^{s})$ to avoid the interference brought by mask contours. Then the scene code $\mathcal{C}_{s}$ and the occlusion code $\mathcal{C}_{o}$ are concatenated together in order to get the full scene code $\mathcal{C}_{so}$ for composed synthesis. The independent encoding of spatial components (i.e., middle human, underlying scene, and floating occlusion) enable the network to learn an automatic layer composition, thus achieving natural character insertion in complicated scenes even with occluded object interactions.
@@
Figure 5: Results of animating diverse characters (e.g., realistic humans, cartoon characters and personified ones) with novel 3D motions retrieved from the motion database (a) or extracted from the driving video (b), and interactive scenes from in-the-wild videos (c).
@@ 3.4 Composed decoding
Given the latent codes of decomposed attributes, we re-compose them as conditions of the diffusion-based decoder for video reconstruction. As shown in Figure 4, we adapt denoising U-Net backbone built upon Stable Diffusion (SD) [34] with temporal layers from [6]. The full scene code $\mathcal{C}_{so}$ is concatenated with the latent noise, and is fed into a 3D convolution layer for fusion and alignment. The motion code $\mathcal{C}_{mo}$ is added to the fused feature and input to the denoising U-Net. For identity code $\mathcal{C}_{id}$, its local feature and global feature are inserted into the U-Net via self-attention layers and cross-attention layers as [12], respectively. Finally, the denoised result is converted into the video clip $\hat{v}$ via a pretrained VAE decoder [20].
@@ 3.5 Training
For the training, we employ the diffusion noise-prediction loss to simulate video reconstruction process:
@@
where $x_{0}$ is the augmented input sample, $t$ denotes the diffusion timestep, $x_{t}$ is the noised sample at $t$, and $\epsilon_{\theta}$ represents the function of the denoising UNet.
Implementation details. Our method is implemented in PyTorch using 8 NVIDIA Tesla-A100 GPUs with 80GB memory. We initialize the model of denoising U-Net and reference-net based on the pre-trained weights from SD 1.5 [34], whereas the motion module is initialized with the weights of AnimateDiff [6]. During training, the weights of VAE encoder and decoder, as well as the CLIP image encoder are frozen. We optimize the denoising U-Net, pose encoder and reference-net with the diffusion noise-prediction loss. It takes around 50k iterations with 24 video frames and a batch size of 8 for converge.
Figure 6: Qualitative comparison with three state-of-the-art methods: Animate Anyone [12], Mimic-Motion [50] and Champ [53].